please help me,
Area of a triangle is 30 square meter. Find the base and altitude, if the altitude exceeds the base by 7 meters. please....

Answer by Alan3354(69443)

Area of a triangle is 30 square meter. Find the base and altitude, if the altitude exceeds the base by 7 meters.
=====================
Area = b*h/2 = 30
b*h = 60
--> 5 & 12